Leonardo, who is married but files separately, earns $90,000 of taxable income. He also has $8,750 in city of Tulsa bonds. His wife, Theresa, earns $37,500 of taxable income.

If Leonardo instead had $27,500 of additional tax deductions for year 2018, his marginal tax rate on the deductions would be: (Use tax rate schedule)

Multiple Choice

  • 12.55%

  • 14.43%

  • 22.55%

  • 16.92%

  • None of the choices are correct.
Answer #1

Answer : None of the choices are correct

Explanation:

Total taxable income of Leonardo = 90000 + 27500 = 117500

Tax rate applicable for the year 2017 for married people filing separately is 28% + 14693.75 for taxable income over $75600.

Marginal rate = (Total tax at $110000 - total tax at $90000)/($110000 - $90000)

= (28%(117500 - 75600) + 14693.75) - 28%(90000 - 75600) + 14693.75)]/(117500 - 90000)

= (15846.25 - 8146.25) / 27500

= 7700/27500

= 0.28 * 100

= 28% Hence none of the choices are correct.